# TPV: Parameter Perturbations Through the Lens of Test Prediction Variance

### v1

@@ Abstract
Theoretically, we show that TPV estimated on the training set converges to its test-set value in the overparameterized limit, providing the first result that prediction variance under local parameter perturbations can be inferred from training inputs alone. Empirically, TPV exhibits a striking stability across datasets and architectures—including extremely narrow networks— and correlates well with clean test loss.
@@ 1 Introduction
Despite remarkable empirical success, our understanding of why modern deep networks generalize—even in heavily overparameterized and noise–perturbed regimes—remains incomplete. Several perspectives have been proposed—wide minima (hochreiter1997flat; keskar2017large), implicit optimization bias (mandt2017stochastic; smith2018bayesian; chaudhari2018stochastic; soudry2018implicit; zhang2021understanding), benign overfitting in overparameterized networks (belkin2019reconciling; bartlett2020benign; belkin2020two), and Neural Tangent Kernel (NTK) (jacot2018neural; adlam2020neural), to name a few. While each viewpoint explains part of the puzzle, they rely on different analytical lenses and seldom yield a single quantity that directly predicts test-set behavior for a fixed, trained model subject to realistic perturbations.
@@
where $H_{\mathrm{eff}}$ is second moment of the output-parameter Jacobian and $C=\mathbb{E}[\delta w\delta w^{\top}]$ is the perturbation covariance. This decomposition separates a label-free geometric factor $H_{\mathrm{eff}}$ from the noise mechanism encoded in $C$ and reveals that diverse perturbations—including SGD noise, label noise, quantization noise, and pruning masks—all influence test predictions through the same curvature–covariance interaction. Variants of $\mathrm{Tr}(H_{\mathrm{eff}}C)$ appear in analyses of SGD dynamics (zhu2018anisotropic; thomas2020interplay; bar2024expected). However, prior work does not interpret this object as a predictive variance functional nor use it as a unifying lens for heterogeneous real-world perturbations.
@@
    TPV Trace Stability: We prove that, in overparameterized networks, training-set TPV converges to test-set TPV, providing the first theoretical result showing that logit prediction variance under parameter perturbations, when evaluated on training-set inputs, is a reliable test-time estimator. Empirically, we find that TPV stability holds true even at very low widths and only breaks if either the number of samples are low or the first order approximation is broken by perturbations.
@@ 2 Test Prediction Variance
Classical bias–variance analyses (geman1992neural; friedman1997biasvariance) study the global prediction variance $\operatorname{Var}_{w}[f_{w}(x)]$ obtained by retraining the model under different sources of randomness (e.g., parameter initialization, stochastic gradient noise, stochastic regularization such as dropout). For instance, neal2018modern revisit this global notion for deep networks, analyzing the variability induced by initialization, data sampling, and optimization. Accordingly, global variance provides insight into properties of the learning algorithm—such as algorithmic stability, the double descent phenomenon, and the benefits of ensembling.
@@ 2.2 TPV Trace Stability
In this section, we prove that for trained overparameterized networks of width $m\rightarrow\infty$, TPV estimated entirely using the training dataset acts as a good estimator of the test set TPV object.
@@
See Appendix A for proof. This result is important because it shows that, in sufficiently overparameterized networks, the TPV estimate $H_{\mathrm{eff}}(w^{\star};X_{\mathrm{tr}})$ and equivalently Eq. 2 (under the first order approximation) estimated using the training set already carries enough information to approximate the test-set TPV that governs test prediction variance.
The intuition behind the proof is simple and uses two observations–1. jacot2018neural; allen2019convergence show that the NTK kernel remains stable during training; 2. due to random initialization and i.i.d. sampling of datasets, $H_{\mathrm{eff}}(w_{0};X)$ concentrates around its population counterpart in operator norm for both $X_{\mathrm{tr}}$ and $X_{\mathrm{te}}$ due to the Law of Large Numbers. Combining these two observations allows us to upper-bound the distance between the training set and test set TPV objects. We use this assumption in section 3.2 and 3.3.
@@ 3 Parameter Perturbation Sources
In this section, we study the effect of individual noise sources in isolation on test prediction variance (TPV). Specifically, we revisit label noise, SGD mini-batch noise, and finite-precision noise, and show how the benefits of over-parameterization and wide minima arise from the common mechanism of suppressing TPV.
@@ 3.1 Label Noise
The proof is shown in Appendix B for completeness. Note that in the linear regression setting above, the “global” prediction variance $\mathbb{E}_{x}\operatorname{Var}_{\varepsilon}(f_{w^{\star}}(x))$ obtained by retraining on different noisy label realizations coincides exactly with the local TPV quantity $\mathrm{Tr}(H_{\mathrm{eff}}C)$, and is a well known result (bartlett2020benign) in literature characterizing benign overfitting in linear regression models. This is because the conditioning of the matrix $XX^{\top}$ improves as $d$ grows beyond $n$, resulting in lower $\operatorname{TPV}_{\mathrm{label}}$, making the model less sensitive to label noise. Even more specifically, if training and test distributions are identical, then for a large enough $n(\ll d)$, we have $\frac{1}{n}\,X^{\top}X~=I_{d}$ (due to the whitening assumption). Under isotropic/whitened assumptions on the rows of $X$, random matrix theory (Wishart) gives $\mathbb{E}(XX^{\top})^{-1}\approx\frac{1}{d}I_{n}$. Therefore, $\operatorname{TPV}_{\mathrm{label}}\approx\sigma_{\varepsilon}^{2}n/d$.
@@
The proof is provided in Appendix C. Equation 10 expresses the contribution of label noise to test prediction variance as a spectral sum involving (i) test-distribution Jacobian modes $B_{ii}$ and (ii) the corresponding training singular values $s_{i}$. The linear result (Eq. 9) is a special case of Eq. 10 when the nonlinear Jacobian is replaced by the data matrix.. To our knowledge, Eq. 10 is the first expression of label-noise–induced prediction variance for finite-width deep networks in terms of a test-Jacobian operator and a general parameter-noise covariance $C$, that also accommodates other perturbation sources (e.g., SGD noise, finite precision, pruning masks) within the same $\operatorname{Tr}(H_{\mathrm{eff}}C)$ template.
@@
Let $J(w)\in\mathbb{R}^{n\times p}$ denote the Jacobian of network outputs on the $n$ training points and let $G(w)=J(w)J(w)^{\top}$ be the empirical NTK matrix. When $p\gg n$, the non-zero singular values of $J(w)$ correspond to the eigenvalues of $G(w)$. For two-layer ReLU networks, du2018gradient show that with sufficient width, the smallest eigenvalue $\lambda_{\min}(G(w_{0}))$ is bounded away from zero at initialization and remains so during training. bombari2022memorization extend this to deep networks with $\Omega(n)$ parameters, establishing well-conditioned NTK structure at initialization. Moreover, allen2019convergence prove that SGD remains in a neighborhood of initialization where the NTK matrix stays close to its well-conditioned initial value.
@@
Two pathological regimes are worth noting w.r.t. Eq. 10. (i) When the linearized system around $w^{\star}$ interpolates the noisy labels (e.g., $J\delta w=\varepsilon$ with full row rank), both the training TPV and the label-noise TPV in Eq. 10 collapse to $\sigma_{\varepsilon}^{2}$ (or to 0 as $\sigma_{\varepsilon}^{2}\!\to\!0$) if the training and test distributions match. TPV stability then holds trivially, but TPV becomes uninformative across architectures. (ii) When $\sigma_{\varepsilon}^{2}$ is sufficiently large that the noisy solution leaves the linearization regime, the assumptions of Theorem 3.2 no longer apply and the label-noise TPV formula should not be used. The non-degenerate, small-but-finite noise regime—where interpolation does not occur and the linear approximation remains valid, and the min-norm solution is achieved—is the regime in which label noise TPV exhibits meaningful geometric dependence and is useful. On a related note, SGD in deep networks may not find the min-norm solution in $\delta w$ and is a proxy for the true optimization problem. See Appendix C.2 and C.3 for a detailed discussion on the label noise TPV object.
@@ 3.2 SGD Stationary Noise Near Convergence
where, $\eta$ and $b$ are the SGD learning rate and batch size, and $\sigma_{\varepsilon}^{2}$ denotes the variance of the residual error over the training samples induced by mini-batch gradient noise.
@@
Connection to Wide-minima Hypothesis: There has been much debate on whether or not the flatness of minima affects the generalization performance of the final model achieved at the end of the training process. The classic counter-argument against the wide-minima hypothesis by dinh2017sharp is that by reparameterizing the weights of a network, the spectrum of the Hessian can be changed without changing the input-output function map.
@@ 3.4 Jacobian-Based Rebalancing (JBR) for Pruning
The proposed JBR is very similar to Jacobian Criterion (JC) proposed by (chen2025optimal) and can be seen as a label-free version of JC. Both JBR and JC assign a score to each parameter group $g$ of the form
@@ 4.1 Test Prediction Variance
Work on double descent and benign overfitting explains how heavily overparameterized models can interpolate noisy data yet still generalize, typically by analyzing the test risk of minimum-norm or ridgeless interpolators as a function of model capacity or feature spectrum (belkin2019reconciling; bartlett2020benign). This line of work focuses on global risk curves and asymptotic regimes, rather than on how small parameter perturbations around a fixed trained model translate into test prediction variance.
@@
Work on implicit optimization bias studies how training dynamics alone can select particular predictors among many interpolating solutions, for example modeling SGD as a stochastic process that samples from a local Gibbs/posterior distribution (mandt2017stochastic; smith2018bayesian; chaudhari2018stochastic) or showing that (stochastic) gradient descent on separable problems converges to max-margin classifiers in function space (soudry2018implicit), or the implicit bias of SGD towards solutions that generalize well (zhang2021understanding; rahaman2019spectral; xu2019frequency). These works focus on characterizing the solutions and stationary distributions implicitly favored by the optimizer. Similarly, classical work on wide/flat minima (hochreiter1997flat; keskar2017large) and related sharpness-aware training methods such as SAM (foret2020sharpness) are concerned with how the geometry of the loss landscape—typically measured through Hessian eigenvalues or sharpness—correlates with generalization. These approaches focus on the selection process of wide versus flat minima and how optimization affects this choice. The TPV framework analyzes how curvature interacts with parameter perturbation covariance for a trained model for arbitrary perturbation sources and links local parameter sensitivity to test-time behavior.
@@
Classical bias–variance analyses decompose prediction error into contributions from the mean predictor and its variability manifested by retraining the model on different training sets and other sources of randomness (geman1992neural; friedman1997biasvariance; neal2018modern). More recently, bordelon2023dynamics analyze the prediction fluctuations in finite-width networks using dynamical mean-field theory to characterize how test prediction variance arises from kernel fluctuations during training. In contrast to these global decompositions of test risk, our TPV framework focuses on a local prediction-variance functional around a fixed trained model, quantifying how curvature and parameter perturbation covariance jointly control test prediction variance.
@@
There is also research on bias-variance decomposition that studies how the two components correlate in modern deep learning. neal2018modern show that, as network width grows, both the global bias and variance components of test error can decrease together, contradicting the textbook U-shaped trade-off. yang2020rethinking similarly find monotonically decreasing bias and a unimodal (bell-shaped) variance curve in width, so that in the overparameterized regime risk is largely controlled by bias. A complementary random-feature study by adlam2020understanding provides a fine-grained global bias–variance decomposition and identifies regimes where both terms become small. In contrast, our TPV formulation investigates a fundamentally different phenomenon: a local, perturbation-based variance (TPV) around a fixed trained reference model, exhibits a systematic correlation with clean test error.
@@
Work on the Neural Tangent Kernel (NTK) shows that, in the infinite-width limit, gradient descent training of neural networks is equivalent to kernel regression with a fixed kernel, yielding precise characterizations of training dynamics and generalization in terms of the NTK spectrum (jacot2018neural; lee2019wide; arora2019exact). Our TPV framework leverages this infinite-width NTK regime to obtain a trace-stability theorem —showing that $\mathrm{Tr}\bigl(H_{\mathrm{eff}}\,C_{\mathrm{prune}}\bigr)$ computed on training set converges to the test set estimate—while providing a curvature–covariance interpretation of prediction variance applicable to finite networks as well.
@@ 4.2 Pruning
Pruning aims at reducing model size and improving inference speed (lecun1989optimal; hassibi1993optimal). These methods can be broadly categorized into unstructured and structured pruning.
Unstructured pruning removes each parameter individually  (han2015learning; wang2020picking; frankle2018lottery; paul2022unmasking). While this approach aligns with the goal of reducing model size, it often does not significantly boost inference speed. Structured pruning (molchanov2019importance; liu2021group), on the other hand, removes entire neurons, filters, or attention heads, resulting in models that are easier to accelerate on standard hardware.
Structured pruning can be further divided into data-dependent and data-independent pruning. Data-independent strategies typically rely on pretrained weight statistics; removing groups with small parameter norm (Li2017Pruning), BatchNorm scale (liu2017learning), or geometric and structural properties of the pretrained weights—such as norms, redundancy, clustering, or subspace contribution—to estimate filter importance without relying on labels or loss gradients (he2019filter; singh2020leveraging; chen2023whc). Data-dependent strategies on the other hand prune parameter groups based on the sensitivity of loss w.r.t. neurons or parameters. The sensitivity can be estimated in different ways– second order approximation (liu2021group), Fisher approximation (theis2018faster), and first order approximation (you2019gate; molchanov2016pruning; molchanov2019importance; chen2025optimal).
The proposed JBR pruning strategy is particularly close to Jacobian Criterion (JC) (chen2025optimal). The key difference is that JC measures loss sensitivity w.r.t. parameter groups using ground-truth labels under the first order approximation, while JBR measures loss sensitivity using the model predicted class labels for confident samples. Under the specific scenario where ground-truth labels are fully correct and model predictions are fully accurate on the training samples, the two become equivalent. The main motivation behind JBR is that it models pruning as a special case of parameter perturbation noise under the general TPV framework.

### v2

@@ Abstract
Theoretically, we show that TPV estimated on the training set converges to its test-set value in the overparameterized limit, providing the first result that prediction variance under local parameter perturbations can be inferred from training inputs alone, and this stability is decoupled from generalization performance. Empirically, TPV exhibits a striking stability across datasets and architectures—including extremely narrow networks— and TPV correlates well with clean test loss.
@@ 1 Introduction
Despite remarkable empirical success, our understanding of why modern deep networks generalize—even in heavily overparameterized and noise–perturbed regimes—remains incomplete. Several perspectives have been proposed—wide minima (Hochreiter and Schmidhuber, 1997; Keskar and et al., 2017), implicit optimization bias (Mandt et al., 2017; Smith and Le, 2018; Chaudhari and Soatto, 2018; Soudry et al., 2018; Zhang et al., 2017), benign overfitting in overparameterized networks (Belkin et al., 2019; Bartlett et al., 2020; Belkin et al., 2020), and Neural Tangent Kernel (NTK) (Jacot et al., 2018; Adlam and Pennington, 2020a), to name a few. While each viewpoint explains part of the puzzle, they rely on different analytical lenses and seldom yield a single quantity that directly predicts test-set behavior for a fixed, trained model subject to realistic perturbations.
@@
where $H_{\mathrm{eff}}$ is second moment of the output-parameter Jacobian and $C=\mathbb{E}[\delta w\delta w^{\top}]$ is the perturbation covariance. This decomposition separates a label-free geometric factor $H_{\mathrm{eff}}$ from the noise mechanism encoded in $C$ and reveals that diverse perturbations—including SGD noise, label noise, quantization noise, and pruning masks—all influence test predictions through the same curvature–covariance interaction. Variants of $\mathrm{Tr}(H_{\mathrm{eff}}C)$ appear in analyses of SGD dynamics (Zhu et al., 2018; Thomas et al., 2020; Bar et al., 2024). However, prior work does not interpret this object as a predictive variance functional nor use it as a unifying lens for heterogeneous real-world perturbations.
@@
    TPV Trace Stability: We prove that, in overparameterized networks, training-set TPV converges to test-set TPV, providing the first theoretical result showing that logit prediction variance under parameter perturbations, when evaluated on training-set inputs, is a reliable test-time estimator, irrespective of the model’s generalization performance. Empirically, we find that TPV stability holds true even at very low widths and only breaks if either the number of samples are low or the induced perturbations are too large.
@@ 2 Test Prediction Variance
Classical bias–variance analyses (Geman et al., 1992; Friedman, 1997) study the global prediction variance $\operatorname{Var}_{w}[f_{w}(x)]$ obtained by retraining the model under different sources of randomness (e.g., parameter initialization, stochastic gradient noise, stochastic regularization such as dropout). For instance, Neal et al. (2018) revisit this global notion for deep networks, analyzing the variability induced by initialization, data sampling, and optimization. Accordingly, global variance provides insight into properties of the learning algorithm—such as algorithmic stability, the double descent phenomenon, and the benefits of ensembling.
@@ 2.2 TPV Trace Stability
In this section, we prove that for trained overparameterized networks of width $m\rightarrow\infty$, TPV estimated entirely using the training dataset acts as a good estimator of the test set TPV object, irrespective of a model’s generalization performance.
@@
See Appendix A for proof. This result is important because it shows that, in sufficiently overparameterized networks, the TPV estimate $\mathrm{Tr}(H_{\mathrm{eff}}(w^{\star};X_{\mathrm{tr}})\,C)$ and equivalently Eq. 2 (under the first order approximation) estimated using the training set already carries enough information to approximate the test-set TPV that governs test prediction variance.
The intuition behind the proof is simple and uses two observations–1. Jacot et al. (2018); Allen-Zhu et al. (2019) show that the NTK kernel remains stable during training; 2. due to random initialization and i.i.d. sampling of datasets, $H_{\mathrm{eff}}(w_{0};X)$ concentrates around its population counterpart in operator norm for both $X_{\mathrm{tr}}$ and $X_{\mathrm{te}}$ due to the Law of Large Numbers. Combining these two observations allows us to upper-bound the distance between the training set and test set TPV objects. We use this assumption in section 3.2 and 3.3.
@@ 3 Parameter Perturbation Sources
In this section, we study the effect of individual noise sources in isolation on test prediction variance (TPV). Specifically, we revisit label noise, SGD mini-batch noise, and finite-precision noise, and pruning noise and show how the benefits of over-parameterization and wide minima arise from the common mechanism of suppressing TPV. We also model pruning as a TPV perturbation and propose an algorithm that performs competitively with the state-of-the-art baselines.
@@ 3.1 Label Noise
The proof is shown in Appendix B for completeness. Note that in the linear regression setting above, the “global” prediction variance $\mathbb{E}_{x}\operatorname{Var}_{\varepsilon}(f_{w^{\star}}(x))$ obtained by retraining on different noisy label realizations coincides exactly with the local TPV quantity $\mathrm{Tr}(H_{\mathrm{eff}}C)$, and is a well known result (Bartlett et al., 2020) in literature characterizing benign overfitting in linear regression models. This is because the conditioning of the matrix $XX^{\top}$ improves as $d$ grows beyond $n$, resulting in lower $\operatorname{TPV}_{\mathrm{label}}$, making the model less sensitive to label noise. Even more specifically, if training and test distributions are identical, then for a large enough $n(\ll d)$, we have $\frac{1}{n}\,X^{\top}X~=I_{d}$ (due to the whitening assumption). Under isotropic/whitened assumptions on the rows of $X$, random matrix theory (Wishart) gives $\mathbb{E}(XX^{\top})^{-1}\approx\frac{1}{d}I_{n}$. Therefore, $\operatorname{TPV}_{\mathrm{label}}\approx\sigma_{\varepsilon}^{2}n/d$.
@@
The proof is provided in Appendix C. Equation 10 expresses the contribution of label noise to test prediction variance as a spectral sum involving (i) test-distribution Jacobian modes $B_{ii}$ and (ii) the corresponding training singular values $s_{i}$. The linear result (Eq. 9) is a special case of Eq. 10 when the nonlinear Jacobian is replaced by the data matrix. To our knowledge, Eq. 10 is the first expression of label-noise–induced prediction variance for finite-width deep networks in terms of a test-Jacobian operator and a general parameter-noise covariance $C$, that also accommodates other perturbation sources (e.g., SGD noise, finite precision, pruning masks) within the same $\operatorname{Tr}(H_{\mathrm{eff}}C)$ template.
@@
Let $J(w)\in\mathbb{R}^{n\times p}$ denote the Jacobian of network outputs on the $n$ training points and let $G(w)=J(w)J(w)^{\top}$ be the empirical NTK matrix. When $p\gg n$, the non-zero singular values of $J(w)$ correspond to the eigenvalues of $G(w)$. For two-layer ReLU networks, Du et al. (2018) show that with sufficient width, the smallest eigenvalue $\lambda_{\min}(G(w_{0}))$ is bounded away from zero at initialization and remains so during training. Bombari et al. (2022) extend this to deep networks with $\Omega(n)$ parameters, establishing well-conditioned NTK structure at initialization. Moreover, Allen-Zhu et al. (2019) prove that SGD remains in a neighborhood of initialization where the NTK matrix stays close to its well-conditioned initial value.
@@
Two pathological regimes are worth noting w.r.t. Eq. 10. (i) When the linearized system around $w^{\star}$ interpolates the noisy labels (e.g., $J\delta w=\varepsilon$ with full row rank), both the training TPV and the label-noise TPV in Eq. 10 collapse to $\sigma_{\varepsilon}^{2}$ (or to 0 as $\sigma_{\varepsilon}^{2}\!\to\!0$) if the training and test distributions match. TPV stability then holds trivially, but TPV becomes uninformative across architectures. (ii) When $\sigma_{\varepsilon}^{2}$ is sufficiently large that the noisy solution either leaves the linearization regime (breaking the assumptions of Theorem 3.2) or induces a large $\mathrm{Tr}(C)$ making the upper bound in Theorem 2.1 weak, it can both cause TPV stability to break. The non-degenerate, small-but-finite noise regime—where interpolation does not occur and sufficiently small perturbations are used, and the min-norm solution is achieved—is the regime in which label noise TPV exhibits meaningful geometric dependence and is useful. On a related note, SGD in deep networks may not find the min-norm solution in $\delta w$ and is a proxy for the true optimization problem. See Appendix C.2 and C.3 for a detailed discussion on the label noise TPV object.
@@ 3.2 SGD Stationary Noise Near Convergence
where, $\eta$ and $b$ are the SGD learning rate and batch size, and $\sigma_{\varepsilon}^{2}$ denotes the variance of the residual error over the training samples.
@@
Connection to Wide-minima Hypothesis: There has been much debate on whether or not the flatness of minima affects the generalization performance of the final model achieved at the end of the training process. The classic counter-argument against the wide-minima hypothesis by Dinh et al. (2017) is that by reparameterizing the weights of a network, the spectrum of the Hessian can be changed without changing the input-output function map.
@@ 3.4 Jacobian-Based Rebalancing (JBR) for Pruning
The proposed JBR is very similar to Jacobian Criterion (JC) proposed by (Chen et al., 2025) and can be seen as a label-free version of JC. Both JBR and JC assign a score to each parameter group $g$ of the form
@@ 4.1 Test Prediction Variance
Work on double descent and benign overfitting explains how heavily overparameterized models can interpolate noisy data yet still generalize, typically by analyzing the test risk of minimum-norm or ridgeless interpolators as a function of model capacity or feature spectrum (Belkin et al., 2019; Bartlett et al., 2020). This line of work focuses on global risk curves and asymptotic regimes, rather than on how small parameter perturbations around a fixed trained model translate into test prediction variance.
@@
Work on implicit optimization bias studies how training dynamics alone can select particular predictors among many interpolating solutions, for example modeling SGD as a stochastic process that samples from a local Gibbs/posterior distribution (Mandt et al., 2017; Smith and Le, 2018; Chaudhari and Soatto, 2018) or showing that (stochastic) gradient descent on separable problems converges to max-margin classifiers in function space (Soudry et al., 2018), or the implicit bias of SGD towards solutions that generalize well (Zhang et al., 2017; Rahaman et al., 2019; Xu et al., 2019). These works focus on characterizing the solutions and stationary distributions implicitly favored by the optimizer. Similarly, classical work on wide/flat minima (Hochreiter and Schmidhuber, 1997; Keskar and et al., 2017) and related sharpness-aware training methods such as SAM (Foret et al., 2020) are concerned with how the geometry of the loss landscape—typically measured through Hessian eigenvalues or sharpness—correlates with generalization. These approaches focus on the selection process of wide versus flat minima and how optimization affects this choice. The TPV framework analyzes how curvature interacts with parameter perturbation covariance for a trained model for arbitrary perturbation sources and links local parameter sensitivity to test-time behavior.
@@
Classical bias–variance analyses decompose prediction error into contributions from the mean predictor and its variability manifested by retraining the model on different training sets and other sources of randomness (Geman et al., 1992; Friedman, 1997; Neal et al., 2018). More recently, Bordelon and Pehlevan (2023) analyze the prediction fluctuations in finite-width networks using dynamical mean-field theory to characterize how test prediction variance arises from kernel fluctuations during training. In contrast to these global decompositions of test risk, our TPV framework focuses on a local prediction-variance functional around a fixed trained model, quantifying how curvature and parameter perturbation covariance jointly control test prediction variance.
@@
There is also research on bias-variance decomposition that studies how the two components correlate in modern deep learning. Neal et al. (2018) show that, as network width grows, both the global bias and variance components of test error can decrease together, contradicting the textbook U-shaped trade-off. Yang et al. (2020) similarly find monotonically decreasing bias and a unimodal (bell-shaped) variance curve in width, so that in the overparameterized regime risk is largely controlled by bias. A complementary random-feature study by Adlam and Pennington (2020b) provides a fine-grained global bias–variance decomposition and identifies regimes where both terms become small. In contrast, our TPV formulation investigates a fundamentally different phenomenon: a local, perturbation-based variance (TPV) around a fixed trained reference model, exhibits a systematic correlation with clean test error.
@@
Work on the Neural Tangent Kernel (NTK) shows that, in the infinite-width limit, gradient descent training of neural networks is equivalent to kernel regression with a fixed kernel, yielding precise characterizations of training dynamics and generalization in terms of the NTK spectrum (Jacot et al., 2018; Lee et al., 2019; Arora et al., 2019). Our TPV framework leverages this infinite-width NTK regime to obtain a trace-stability theorem —showing that $\mathrm{Tr}\bigl(H_{\mathrm{eff}}\,C_{\mathrm{prune}}\bigr)$ computed on training set converges to the test set estimate—while providing a curvature–covariance interpretation of prediction variance applicable to finite networks as well.
@@ 4.2 Pruning
Pruning aims at reducing model size and improving inference speed (LeCun et al., 1989; Hassibi et al., 1993). These methods can be broadly categorized into unstructured and structured pruning.
Unstructured pruning removes each parameter individually  (Han et al., 2015; Wang et al., 2020; Frankle and Carbin, 2018; Paul et al., 2022). While this approach aligns with the goal of reducing model size, it often does not significantly boost inference speed. Structured pruning (Molchanov et al., 2019; Liu et al., 2021), on the other hand, removes entire neurons, filters, or attention heads, resulting in models that are easier to accelerate on standard hardware.
Structured pruning can be further divided into data-dependent and data-independent pruning. Data-independent strategies typically rely on pretrained weight statistics; removing groups with small parameter norm (Li et al., 2017), BatchNorm scale (Liu et al., 2017), or geometric and structural properties of the pretrained weights—such as norms, redundancy, clustering, or subspace contribution—to estimate filter importance without relying on labels or loss gradients (He et al., 2019; Singh et al., 2020; Chen et al., 2023). Data-dependent strategies on the other hand prune parameter groups based on the sensitivity of loss w.r.t. neurons or parameters. The sensitivity can be estimated in different ways– second order approximation (Liu et al., 2021), Fisher approximation (Theis et al., 2018), and first order approximation (You et al., 2019; Molchanov et al., 2016, 2019; Chen et al., 2025).
The proposed JBR pruning strategy is particularly close to Jacobian Criterion (JC) (Chen et al., 2025). The key difference is that JC measures loss sensitivity w.r.t. parameter groups using ground-truth labels under the first order approximation, while JBR measures loss sensitivity using the model predicted class labels for confident samples. Under the specific scenario where ground-truth labels are fully correct and model predictions are fully accurate on the training samples, the two become equivalent. The main motivation behind JBR is that it models pruning as a special case of parameter perturbation noise under the general TPV framework.